# Quaternion wavefunction formulation of incompressible inviscid fluid dynamics

### v3

@@ Appendix B Quaternion Logarithm Gradient
    $=\frac{\nabla\rho}{2\rho}+\frac{i\nabla S}{\hbar_{f}}\cdot\bm{\sigma}+\frac{\ell}{\hbar_{f}}\nabla_{Q}\star(\mathbf{v}\cdot\bm{\sigma}).$ (50)
@@
    $\nabla_{Q}\star(\mathbf{v}\cdot\bm{\sigma})=(\nabla\cdot\mathbf{v})+(\nabla\mathbf{v})\cdot\bm{\sigma}.$ (51)
@@
    $\mathrm{Re}[\nabla_{Q}\star\ln\Psi]=\frac{\nabla\cdot(\nabla\rho)}{2\rho}+\frac{\ell}{\hbar_{f}}(\nabla\cdot\mathbf{v}).$ (52)
    For constant density $\rho=\rho_{0}$:
@@
    Appendix C Recovery of Euler Equations from GPE
    We derive Eqs. (20a)–(20c) from the GPE. Starting with $\Psi=\sqrt{\rho}\exp(iS/\hbar_{f})q_{v}$ and $\partial_{t}\Psi$:
@@ Appendix C Recovery of Euler Equations from GPE
    Real scalar part:
@@
    Imaginary scalar part:
@@
    Quaternion vector part:

### v4

@@ Appendix C Quaternion logarithm gradient
$=\frac{\nabla\rho}{2\rho}\cdot\bm{\sigma}+\frac{i\nabla S}{\hbar_{f}}\cdot\bm{\sigma}+\frac{\ell}{\hbar_{f}}\nabla_{Q}\star(\mathbf{v}\cdot\bm{\sigma}).$ (69)
@@
$\nabla_{Q}\star(\mathbf{v}\cdot\bm{\sigma})=-(\nabla\cdot\mathbf{v})+(\nabla\mathbf{v})\cdot\bm{\sigma}.$ (70)
@@
$\mathrm{Re}[\nabla_{Q}\star\ln\Psi]=-\frac{\ell}{\hbar_{f}}(\nabla\cdot\mathbf{v}).$ (71)
For constant density $\rho=\rho_{0}$, the density gradient term vanishes, leaving:
@@
Appendix D Recovery of Euler equations from GPE
We derive Eqs. (20a) through (20c) from the GPE. Starting with $\Psi=\sqrt{\rho}\exp(iS/\hbar_{f})q_{v}$ and computing $\partial_{t}\Psi$:
@@ Appendix D Recovery of Euler equations from GPE
For the real scalar part:
@@
For the imaginary scalar part:
@@
For the quaternion vector part:
@@
With $\Psi=\sqrt{\rho}e^{iS/\hbar_{f}}q_{v}$ and constant $\rho=\rho_{0}$, substitution into the GPE (19) and quaternion decomposition yields:
For the scalar part: $\partial_{t}S+(\nabla S)^{2}/2+\lambda+g\rho_{0}=0$
For the vector part: $\partial_{t}\mathbf{v}+(\mathbf{v}\cdot\nabla)\mathbf{v}=-\nabla[(\lambda+g\rho_{0})/\rho_{0}]$
For the constraint: $\nabla\cdot\mathbf{v}=0$
Identifying $p=(\lambda+g\rho_{0})/\rho_{0}$ recovers Euler equations exactly.
Appendix E Detailed derivation of geometric factors
E.1 Forward factor $\gamma_{Q}^{\text{forward}}=4/9$
The quaternion orthogonality constraint (33) in the attached flow region requires:
$\int_{\Omega_{\text{front}}}(Q\star\nabla_{Q}Q)\cdot(Q\star\nabla_{Q}\bar{Q})^{*}\,dV=0.$ (78)
For potential flow modified by quaternion structure, this constraint determines a reduction factor. Writing the constrained pressure as $C_{p}=\gamma_{Q}C_{p}^{\text{pot}}$, the orthogonality condition becomes:
$\gamma_{Q}^{2}\int C_{p}^{\text{pot}}\cos\theta\sin\theta\,d\theta=\gamma_{Q}\int C_{p}^{\text{pot}}\cos\theta\sin\theta\,d\theta.$ (79)
The non-trivial solution for SU(2) geometry gives $\gamma_{Q}=\cos^{2}\phi$ where $\cos\phi=2/3$, yielding $\gamma_{Q}^{\text{forward}}=4/9$.
E.2 Rear factor $\gamma_{Q}^{\text{rear}}=16/25$
SU(2) complementarity gives $\gamma_{Q}^{\text{base}}=\sin^{2}\phi=5/9$. The wake expansion factor $\eta_{\text{wake}}=(1+\cos\theta_{s})/(1-\cos\theta_{s})$ modifies this through quaternion curvature scaling:
$\gamma_{Q}^{\text{rear}}=\gamma_{Q}^{\text{base}}\cdot\sqrt{\eta_{\text{wake}}}=\frac{5}{9}\cdot 1.19\approx\frac{16}{25}.$ (80)